OpenStudy (anonymous):

The circle given by x^2+y^2-2y-11=0 can be written in standard form like this: x^2+(y-k)^2=12 . What is the value of k in this equation?

OpenStudy (anonymous):

x² + (y² - 2y + ??) = + 11 (1/2)*(-2) = (-1) ; (-1)² = 1......Add 1 to both sides x² + (y² -2y + 1) = 11 + 1 x² + (y - 1)² = 12 ANSWER....k = 1 If you have questions on this or any other math problem feel free to let me know how to contact you.......Read Below....... I tutor free online since moving to France from Florida in June 2011. Hope this is most helpful and you will continue to allow me to tutor you. If you dont fully understand or have a question PLEASE let me know how to contact you so you can then contact me directly with any math questions. If you need me for future tutoring I am not allowed to give out my info but what you tell me on ***how to contact you*** is up to you Been a pleasure to serve you Please call again Robert Jones.............f "Teacher/Tutor of Fine Students"